A box contains 3 red balls and 4 green balls. If two balls are randomly selected in sequence, what is the possibility that a red ball and a green ball are picked out of the box. State the rule for P(A&B) and then substitute the values and compute the answer.

Answer by stanbon(75887)

A box contains 3 red balls and 4 green balls. If two balls are randomly selected in sequence, what is the possibility that a red ball and a green ball are picked out of the box.
----
# of red/green pairs: 3*4 = 12
# of random pairs: 7C2 = (7*6)/(1*2) = 21
---
P(red/greed pair) = 12/21 = 4/7
====================================
Cheers,
Stan H.